OpenStudy (anonymous):

invests $5,000 in an account that pays 6.25% interest per year. After how many years will her investment be worth $10,000?

OpenStudy (lgbasallote):

sounds like a simple interest problem the formula for simple interest is I = Prt in this case P is $5,000; r is 0.0625; I is $10,000 and you're looking for t so just substitute \[\implies 10,000 = (5,000)(0.0625)t\] divide both sides by 5,000 \[\implies 2 = 0.0625 t\] divide both sides by 0.0625 \[\implies 32 = t\] does that help?

OpenStudy (anonymous):

i got 13.7?

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (lgbasallote):

i showed you how to do it. where did you go wrong?

OpenStudy (mimi_x3):

Oh no no..i there's a mistake.. IIRC this is the formula: \[A_n = P\left(1+\frac{r}{100}\right)^{n} \] where r(rate) = 6.25 , n= n , P = 5000, A_n = 10 000

OpenStudy (mimi_x3):

so you're finding n which is the year

OpenStudy (anonymous):

that's for compound interest @Mimi_x3

OpenStudy (mimi_x3):

Compound interest is found by \(A_n - P\)

OpenStudy (mimi_x3):

well i havent done these in a while tho :/

OpenStudy (anonymous):

it's okay @Mimi_x3 page 222 and here's a little chapter http://finedrafts.com/files/Larson%20PreCal%208th/Larson%20Precal%20CH3.pdf

OpenStudy (mimi_x3):

hmm..looks like i have to look back at my notes.. how can you tell the differennce if its simple interest and compound interest? :/

OpenStudy (anonymous):

simple interest only allows the principal to accrue by the term, that is the n year(s) or month(s) compound interest is better understood this way: let us say you are given a choice to pick between 1 million dollars or 1 cent and is doubled every day for the next 30 days, which one would you pick?

OpenStudy (mimi_x3):

oohh..i got it..because the 5k does not double up or anyother words you only invest 5k..i made a mistake then sorry

OpenStudy (anonymous):

here's why banks can rip you off http://www.al6400.com/blog/2006/07/10/a-penny-doubled-everyday/

OpenStudy (anonymous):

14.1??